OpenStudy (goformit100):

A bird shooter was askgd how many birds he had in the bag. He replied that there were all sparrows but six, all pigeons but six, and all ducks but six. How many birds he had in the bag in all? A. 9 B. 18 C. 27 D. 36

OpenStudy (goformit100):

@Eyad Sir I am not knowing its conception

OpenStudy (anonymous):

let total number of birds = x. let no. of sparros = s, pegions = p, ducks = d. given s = x-6. d = x-6. p = x-6. s+d+p = x = 3x-18. so x = 9. there are 9 birds

OpenStudy (goformit100):

First of all i would like to say that Sir you explain very nicely :)